OpenStudy (anonymous):

express 4^-(1/2) = 0.5 in logarithmic form

OpenStudy (turingtest):

\[\log_ax=y\iff x=a^y\]\[a^y=x\iff y=\log_ax\]

OpenStudy (anonymous):

4^(-1/2) = 0.5 take log base 4 both sides \[\log_{4}4^{-1/2}= \log_{4}0.5\]